Let S={(1,4),(0,3)} be a subset of R^2(R).prove that (2,3) belongs to L(S)
1
Expert's answer
2020-12-01T17:29:12-0500

Let's v1 = (1,4) and v2 = (0,3). Now find a and b such that


(2,3) = a*v1 + b*v2

(2,3) = a*(1,4) + b*(0,3)


From here we have linear system

"\begin{cases}
 a*1 + b*0 = 2\\
 a*4 + b*3 = 3
\end{cases}"


From first equation we find a = 2 and substitute it to second we find b = -5/3. Exist a, b in R such that (2,3) = a*v1 + b*v2, where v1, v2 in S


Hence (2,3) belongs to L(S) - the linear span of S.